From GlueXWiki
Jump to: navigation, search

June 14, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production status
    • Spare wire frames (Lubomir)
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Cable assignment and labeling (Vlad)
  5. Gas system (Beni)
  6. Test set-up in 126 (Beni)
  7. Other FDC E-log


Participants: Caleb, Dave, Chris, Nick, Vlad, Simon, Beni, and Lubomir.


- Dave: on last wire frame (#18) had to replace one sense wire and twice one field wire due to incorrect tension, now it's out of the stringing table and Chris will put the components tomorrow. Two wire frames will be deadened tomorrow, and #18 will be deadened on Monday. Cathode #55, the last for package #4, is ready too. On Tuesday/Wednesday we will continue with the stacking of the fourth package. Working on spare cathodes #54, 57, and 59.

- The problem with the noise on cell #4 was solved: a tiny piece of epoxy (or something like that) got stuck to the sense wire in that area. Casey cleaned it but that was not easy. The package was closed on Friday and this week was tested: the noise disappeared.

- Wire frame #1 can't be used at > 1900/-500V and at this voltage the current is very high (~1uA). Before the meeting Eugene agreed to purchase another 4 spare sets of PCBs. A.s.a.p. before purchasing we want to test wire plane #2. It has a conductive spot from the chemical used for deadening, shortening a field and sense wire. We will test also the other two spare wire planes, but they seam to be fine. Then we will know how many new wire planes we need to build. We have only one extra G10 wire frame and in case we need more we have to order them, as well.

Engineering update

- Caleb: Bill is busy, but he explained well details of the outer (outside of the FDC) part of the cooling system that will have Fluorinert and water loops. Caleb will work with Bill also on the testing of the cooling system.

Electronics update

- Chris: LV cables have come. We discussed that it is worth setting up the LV system in 126 and later connecting the LV cables to the third package (when it will be in 126) but not applying voltage on most of them since we will need cooling.

Cable assignment and labeling

- See the document above prepared by Vlad. It contains (in Excel format) a table of all the signal cables showing the label and the destination rack, crate, module, and connector. There's a scheme at the end of the document: half of the cables will go to the left and half to the right racks depending where the corresponding card is. Wire cables will be connected to F1TDCs, 6 modules per package per side occupying two crates with 12 modules each. 5 crates on each side will be filled with 15/14 fADC125s, having 18 modules per package per side. We decided to keep the channel information in the label (like FDC1,C5,D85B-108B).

- The ordering of the connectors is generally (from left to right in the crates) starting from the lower package, cell number, channel number. Beni suggested that we may want to split the cards with highest rates into different fADC modules. It's easy to implement, but we want to discuss this with Fernando, Cody, Gerard.

Gas system

- Beni is in contact with Narciso; soon we will have separate meeting about the gas system.

Test set-up at 126

- Beni: CODA-Light doesn't work with Linux ROCs so far. Will talk to Bryan Moffit if something can be done, if not possible we will have to move to the "heavy" CODA.

- Vlad tested the cathodes of the top three cells of the first package in 126 with a source. The main complain is that on some of the channels there are no signals, but after unscrewing the card from the cooling tube it starts working. It is not clear why: probably the tube pulls/bends the card or the cathode frame. There were suggestions to tighten the tubes to the package evenly around the periphery with cables ties using the holes on the cathodes. Will be discussed with Bill.


- Lubomir looked at some old runs with the second production chamber where we have used also 55Fe source and triggered with one strip. We don't have long tails in the signals (including with comics) as seen in the CDC with fADC125.